Oh, please help me with this, I do NOT understand this system math....
Solve the following system by substitution.
4x - 2y = 8
2x + 4y =8

Answer by vleith(2983)

see this --> http://collabvsl.wetpaint.com/page/Solving+Systems+of+Equations

4x - 2y = 8
2x + 4y =8
2x + 4y = 8
2x = 8 -4y
x = (8-4y)/2 use this and sub into the first equation
4x - 2y = 8
4((8-4y)/2) - 2y = 8
16 - 8y - 2y = 8
-10y = -8
y = 4/5
Use this value of y to sub back in and solve for x
x = 12/5
check your answers